9/6/99 Day Shift Report

KTeV Dart (da_ktev@ktev3)
Mon, 6 Sep 1999 15:37:56 -0500 (CDT)

Shift Summary for day shift, 6/Sep/99

On Shift: VP,RP

Hours of beam available at all intensities: 3h20min
Hours of beam available at nominal intensity: 3h20min
Hours of E832 data (special runs not included): 3h15min
Hours of special runs (laser scans, muon runs...): 1h30min

Accounting of lost time:

========================================================================
Runs ended this shift:

Tapes brought to FCC this shift:
(including tapes for online split)

========================================================================
Accesses during shift: One for changing laser filter wheel position
(There was no beam)

========================================================================
Overview of shift:
E.g., Lots of Beam Interruptions, Change of Laser Filter wheel position,
SDAQ problems (datataking not affected)

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
7.00 Hogan and Jason come out of access. Problems with
large rms errors fixed by another access to reset
filter wheel position 4,0. Jason leaves.

8.30 Problems starting the run - log out/into Ktev3. Reboot
pipeline crate 9. Hogan leaves.

14444 9.10 Start laser scan. 2 3/4 pass scan before abort for
arrival of beam.

14445 9.50 Beam arrives.
9.54 Beam leaves. Left bend ground fault - affects Meson.
10.05 Beam returns.

10.20 NM2S3 trips. MCR requests we ramp down analysis magnet.
No good data so far this run.

14446 10.55 Decide to do laser scan.

11.30 MCR calls to ramp up analysis magnet. Stop run.
(2 3/4 passes once again)

14447 11.40 Beam arrives.

13.43 Data taking proceeding smoothly thus far.

14.05 TRD 6 out of range alarm. Reset the processes. alarm
went away

spill 109 14.15 SDAQ red buttons indicating problems with processes on ATEV2. Val restarts the processes but red persists.
Page Askan who will come in.

14.55 MCR calls to say beam will disappear in about 30 mins
to allow fix left bends ground fault.

15.05 no beam. Ashkan restarts all SDAQ processes
successfully.


========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift